OpenStudy (anonymous):

A ball is tossed into the air from a window. Its path is described by y=-16x^2+8x+12 where y represents the height of the ball in feet above the ground x seconds after it is tossed. what is the vertex form of the equation? what is the maximum height of the ball?

OpenStudy (anonymous):

vertex form: \[f(x) = a(x - h)^2 + k\] where (h, k) is the vertex; meaning k is the max height. you'll want to complete the square: \[y = -16(x^2 - x/2 + 3/4) = -16[(x - 1/4)^2 + 1/4] = -16(x - 1/4)^2 - 4\]

OpenStudy (anonymous):

somethign is wrong with my answer

OpenStudy (anonymous):

i found the mistake

OpenStudy (anonymous):

i wrote +3/4 instead of -3/4 inside the bracket. let me know if you would like me to explain any of my steps ^_^

OpenStudy (anonymous):

\[y = -16(x^2 - x/2 - 3/4) = -16[(x - 1/4)^2 - 11/16] = -16(x - 1/4)^2 + 11\]

OpenStudy (anonymous):

@Euler271 could you explain how you did it please? This all just looks like a big confusing mess to me, :/

OpenStudy (anonymous):

@ehuman @Easyaspi314

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):